OpenStudy (anonymous):

An evergreen nursery usually sells a certain shrub after 6 years of growth and shaping. The growth rate during those 6 years is approximated by dh/dt=3.5t+7 where t is the time in years and h is the height in centimeters. The seedlings are 15 centimeters tall when planted (t=0) . Find the height after t years.

OpenStudy (anonymous):

dh=(3.5t+7)dt Integrating both sides, h=1.75t^2+7t+c where c is an arbitrary constant. You can solve it further from here. t=0, c=15. h=1.74t^2+7t+15

OpenStudy (anonymous):

how do you integrate? @NeetziD

OpenStudy (anonymous):

\[\int\limits_{ }^{} \]

OpenStudy (anonymous):

The RHS is \[\int\limits_{}^{} (3.5t+7) dt\]

OpenStudy (anonymous):

i got h=3.5dt+7dt and thats not an option in my choices